# labelsep = period

### v1

@@ Abstract
Matrix–vector algorithms, particularly Krylov subspace methods, are widely viewed as the most effective algorithms for solving large systems of linear equations. This paper establishes lower bounds on the worst-case number of matrix–vector products needed by such an algorithm to approximately solve a general linear system. The first main result is that, for a matrix–vector algorithm which can perform products with both a matrix and its transpose, $\Omega(\kappa\log(1/\varepsilon))$ matrix–vector products are necessary to solve a linear system with condition number $\kappa$ to accuracy $\varepsilon$, matching an upper bound for conjugate gradient on the normal equations. The second main result is that one-sided algorithms, which lack access to the transpose, must use $n$ matrix–vector products to solve an $n\times n$ linear system, even when the problem is perfectly conditioned. Both main results include explicit constants that match known upper bounds up to a factor of four. These results rigorously demonstrate the limitations of matrix–vector algorithms and confirm the optimality of widely used Krylov subspace algorithms.
@@ 1 Introduction
Solving systems of linear algebraic equations is a fundamental problem in computer science and mathematics, and algorithms for this task are used in almost every area of modern computation, including machine learning [31] and optimization [7]. Given an invertible matrix $\bm{A}\in{\mathbb{R}}^{n\times n}$ and a vector $\bm{b}\in{\mathbb{R}}^{n}$, the objective is to find the vector $\bm{x}\in{\mathbb{R}}^{n}$ for which $\bm{A}\bm{x}=\bm{b}$. The standard algorithms for this task are variants of Gaussian elimination, which run in ${\mathcal{O}}(n^{\omega})$ operations when implemented using fast matrix multiplication. [^1] Here, $\omega\leq 2.371\ldots$ is the matrix multiplication exponent. Determining the optimal complexity for the linear system problem is a major open question [41].
@@
To specify this problem completely, we must specify an accuracy requirement for an approximate solution $\tilde{}\bm{x}$. For this article, we seek a vector $\tilde{}\bm{x}$ with a small relative residual:
@@ 1.1 Background and research questions
Most work on matrix–vector complexity for the $\bm{A}\bm{x}=\bm{b}$ problem focuses on the case when $\bm{A}$ is symmetric positive definite (SPD). In this case, the standard Krylov subspace algorithms are conjugate gradient and MINRES [20, Algs. 2 & 4]. The latter achieves the guarantee ˜1.1 using ${\mathcal{O}}(\sqrt{\kappa}\,\log(\nicefrac{{1}}{{\varepsilon}}))$ matvecs. Here, $\kappa\coloneqq\operatorname{cond}(\bm{A})\coloneqq\lVert\bm{A}\rVert_{2}\lVert\smash{\bm{A}^{-1}}\rVert_{2}=\sigma_{\mathrm{max}}(\bm{A})/\sigma_{\mathrm{min}}(\bm{A})$ is the condition number. Existing lower bounds confirm that $\Omega(\sqrt{\kappa}\,\log(\nicefrac{{1}}{{\varepsilon}}))$ matvecs are necessary for any deterministic algorithm to achieve this guarantee [34, Sec. 7.2]; for randomized algorithms, only a weaker $\Omega(\sqrt{\kappa}/\operatorname{polylog}(\kappa))$ lower bound is known [5]. [^2] If one specializes our arguments to the case of SPD matrices, our techniques yield a $\Omega(\sqrt{\kappa}\log(\nicefrac{{1}}{{\varepsilon}}))$ matvec lower bound against randomized methods; see Section E.2. It is natural to inquire how the complexity changes when we consider general linear systems for which $\bm{A}$ is not SPD.
When considering matrix–vector algorithms for general linear systems, we can distinguish between a two-sided matrix–vector algorithm, which can perform matrix–vector products with both $\bm{A}$ and $\bm{A}^{\smash{\intercal}}$, and a one-sided matrix–vector algorithm, which can only perform matrix–vector products with $\bm{A}$. Both one- and two-sided algorithms are studied and used in practice [20].
@@
The matrix $\bm{A}^{\smash{\intercal}}\bm{A}$ is SPD, so we can solve this system by applying the conjugate gradient algorithm. The condition number of the normal equations is larger $\operatorname{cond}(\bm{A}^{\smash{\intercal}}\bm{A})=\kappa^{2}$, which implies that conjugate gradient on the normal equations requires ${\mathcal{O}}(\sqrt{\kappa^{2}}\kern 0.50003pt\log(\nicefrac{{1}}{{\varepsilon}}))={\mathcal{O}}(\kappa\log(\nicefrac{{1}}{{\varepsilon}}))$ matvecs to produce an $\varepsilon$-accurate solution [33, Eqn. 2.12]. We ask:
@@
The most popular one-sided matrix–vector algorithm is the GMRES algorithm [42]. This algorithm has many desirable properties: When it is applied to an SPD system, it achieves the same ${\mathcal{O}}(\sqrt{\kappa}\kern 0.80002pt\log(\nicefrac{{1}}{{\varepsilon}}))$ complexity as MINRES, and it converges rapidly on many problems in practice. However, this algorithm also has a fundamental weakness: There exists a perfectly conditioned linear system ($\kappa=1$) on which GMRES requires a full $n$ iterations to solve the system to any accuracy level $\varepsilon<1$ [33, Ex. C]. This raises the question:
@@ 1.2 Our results
Fix $\eta>0$. There does not exist any algorithm that takes inputs $\kappa$, $\varepsilon$, and $\bm{b}$, which computes fewer than $\frac{1-\eta}{4}\kappa\log(\nicefrac{{1}}{{\varepsilon}})$ two-sided matrix–vector products with $\bm{A}$, and which returns a vector $\tilde{}\bm{x}$ such that
@@
This result show that conjugate gradient on the normal equations is optimal from the perspective of worst-case computational complexity. [^3] We note that there are reasons in practice to avoid iterative methods based on the normal equations, including numerical stability [35, Sec. 8.2] and better-than-worst case performance [33]. A strength of this result is its explicitness, with the complexity bound featuring no hidden constants. Conjugate gradient on the normal equations achieves the accuracy guarantee ˜1.1 in $(1+o(1))\kappa\log(\nicefrac{{1}}{{\varepsilon}})$ matvecs [20, Eqn. 3.6], so the lower and upper bounds match to a factor of roughly 4. We are unaware of any pair of upper and lower bounds in the matrix–vector complexity literature that come close to this level of parity.
@@
Consider any algorithm that takes a vector $\bm{b}$ and a one-sided matrix–vector oracle for a matrix $\bm{A}$ and outputs a vector $\tilde{}\bm{x}$, and suppose $n$ is larger than some universal constant. Then …
$\text{Achieving }\lVert\bm{A}\tilde{}\bm{x}-\bm{b}\rVert_{2}$ $\leq\frac{1}{2}\lVert\bm{b}\rVert_{2}$ $\text{with prob.\ }\geq\frac{2}{3}$ $\text{requires at least $\lceil n/2\rceil$ matvecs}.$
@@
Algorithms like conjugate gradient and GMRES are already known to be optimal among certain classes of algorithms, like within the class of deterministic single-vector Krylov methods (see [20, Ch. 2] and [34, Sec. 7.2] ). However, prior works do not fully answer if these methods are optimal within the larger class of general randomized matrix–vector algorithms. [^4] Though Braverman et al. [5] do prove that $\Omega(\sqrt{\kappa}/\operatorname{polylog}(\kappa))$ matvecs are needed to solve a SPD linear system to constant error, partially resolving this question. Indeed, there have been many examples of randomized algorithms that achieve surprising improvements to the complexity of problems in linear algebra [39, 10, 11, 32, 29, 36, 15]. However, despite all efforts, no improvements to the worst-case complexity of solving general linear systems have materialized. Our research demonstrates why this is the case: Linear system solving is not vulnerable to such a speedup, in both the one- and two-sided cases.
@@
We use different techniques to handle the one- and two-sided cases. The two-sided lower bound (Theorem˜1.1) uses an approach based on the nonexistence of good polynomial approximations, extending an argument of Chewi et al. [9] . We show that since no polynomial can approximate the function $f(x)=\nicefrac{{1}}{{x}}$ to error $\varepsilon$ on the split interval ${\mathcal{S}}\;{\vcentcolon=}\;[-\kappa,-1]\cup[1,\kappa]$ with degree less that $t={\mathcal{O}}(\kappa\log(1/\varepsilon))$, no algorithm can use fewer than ${\mathcal{O}}(t)$ matvecs to estimate $\operatorname{tr}(f(\bm{A}))=\operatorname{tr}(\bm{A}^{-1})$ for all matrices with eigenvalues in $\mathcal{S}$}. We then make a reduction, showing that any linear system solver satisfying ˜1.1 can be used to estimate $\operatorname{tr}(\bm{A}^{-1})$, resulting in a proof of Theorem˜1.1. See Section˜2 for details.
@@ 1.3 Implication: Fine-grained lower bounds and complexity separation
An important and well-studied example of this arises when $\bm{A}$ has a number of large outlying singular values that blow up its condition number, which arises in smoothed analysis [43], optimization [7], and machine learning [45]. In this setting, conjugate gradient on the normal equations is known to solve such linear systems in an improved ${\mathcal{O}}(k+\kappa_{k}\log(1/\varepsilon))$ matvecs [2], where $\kappa_{k}\coloneqq\sigma_{k+1}(\bm{A})/\sigma_{\rm min}(\bm{A})$ denotes the condition number without its top-$k$ singular values. This more fine-grained notation of complexity, parameterized by both $k$ and $\kappa_{k}$, circumvents our basic two-sided lower bound (Theorem˜1.1). However, our results can be extended to this different complexity measure, again showing that conjugate gradient on normal equations is optimal:
@@
There does not exist an algorithm that takes inputs $k$, $\kappa_{k}$, $\varepsilon$, and $\bm{b}$, which computes fewer than $\Omega(k+\kappa_{k}\log(1/\varepsilon))$ two-sided matrix-vector products with $\bm{A}$ and returns $\tilde{}\bm{x}$ such that
@@
Notably, this result implies a complexity separation between matrix–vector algorithms and general algorithms that can access the entries of $\bm{A}$ directly. If $\bm{A}$ is a dense matrix and we implement matvecs using the standard algorithm, then Theorem˜1.3 shows that $\Omega(n^{2}k+n^{2}\kappa_{k}\log(1/\varepsilon))$ time is needed to solve a linear system. However, recent work [17, 15] shows that if $\bm{A}$ is stored as a dense array, we can solve $\bm{A}\bm{x}=\bm{b}$ in $\widetilde{}{\mathcal{O}}(k^{3}+n^{2}\kappa_{k}\log(1/\varepsilon))$ time. So, for large $n$ and $k$, having access to the matrix offers a direct improvement over matrix–vector access, demonstrating a complexity separation between matrix–vector algorithms and general algorithms. This is the first result of this kind for general, dense linear systems.
@@ 1.4 Related work
Understanding the behavior of Krylov linear solvers has been a decades-long research challenge in numerical analysis. The literature is extensive; see Greenbaum’s book [20] for a summary of theory. Lower bounds for the rate of convergence of Krylov methods can be derived from potential theory [16]. These lower bounds apply only to specific Krylov methods, not general randomized matrix–vector algorithms.
@@
Our work fits into the growing research topic of transpose-free linear algebra [4, 21], which aims to understand the relative power of one- and two-sided algorithms for various linear algebra problems. The results most similar to ours appear in the thesis of Halikias [21, Sec. 5.2], who analyzes one-sided matrix–vector algorithms for solving linear systems and least-squares problems. Her results show that, even in the favorable situation when $\operatorname{cond}(\bm{A})=1$ and we perform n1n-1 matvecs $\bm{A}\bm{z}_{1},\ldots,\bm{A}\bm{z}_{n-1}$, [^5] Halikias also includes as an assumption that the exact solution $\bm{A}^{-1}\bm{b}$ is not in the span of the queries $\{\bm{z}_{1},\ldots,\bm{z}_{n-1}\}$. the data collected by the algorithm is not sufficient to characterize the solution to a system of linear equations: There exists another well-conditioned matrix $\tilde{}\bm{A}$ that is undistinguished by the queries
@@
Our $\Omega(\kappa)$ lower bound for two-sided matrix–vector algorithms mirrors the $\Omega(\kappa)$ lower bound for the quantum linear systems problem [22, p. 150502-3]. In that context, solving implicit sparse systems in ${\mathcal{O}}(\kappa^{1-\eta}\operatorname{polylog}(n))$ time would imply $\mathsf{BQP}=\mathsf{PSPACE}$ and allow for sub-linear simulation of arbitrary quantum circuits. While the $\Omega(\kappa)$ scaling in both settings is striking, we do not know any way to derive one result from the other or reduce them both to some common principle.

### v2

@@ Abstract
Matrix–vector algorithms, particularly Krylov subspace methods, are widely viewed as the most effective algorithms for solving large systems of linear equations. This paper establishes lower bounds on the worst-case number of matrix–vector products needed by such an algorithm to approximately solve a general linear system. The first main result is that, for any matrix–vector algorithm which is allowed the use of randomization and can perform products with both a matrix and its transpose, $\Omega(\kappa\log(1/\varepsilon))$ matrix–vector products are necessary to solve a linear system with condition number $\kappa$ to accuracy $\varepsilon$, matching an upper bound for conjugate gradient on the normal equations. The second main result is that one-sided algorithms, which lack access to the transpose, must use $n$ matrix–vector products to solve an $n\times n$ linear system, even when the problem is perfectly conditioned. Both main results include explicit constants that match known upper bounds up to a factor of four. These results rigorously demonstrate the limitations of matrix–vector algorithms and confirm the optimality of widely used Krylov subspace algorithms.
@@ 1 Introduction
Solving systems of linear algebraic equations is a fundamental problem in computer science and mathematics, and algorithms for this task are used in almost every area of modern computation, including machine learning [34], scientific computing [22], and optimization [7]. Given an invertible matrix $\bm{A}\in{\mathbb{R}}^{n\times n}$ and a vector $\bm{b}\in{\mathbb{R}}^{n}$, the objective is to find the vector $\bm{x}\in{\mathbb{R}}^{n}$ for which $\bm{A}\bm{x}=\bm{b}$. The standard algorithms for this task are variants of Gaussian elimination, which run in ${\mathcal{O}}(n^{\omega})$ operations when implemented using fast matrix multiplication. [^1] Here, $\omega\leq 2.371\ldots$ is the matrix multiplication exponent. Determining the optimal complexity for the linear system problem is a major open question [45].
@@
To specify this problem completely, we must determine an accuracy requirement for an approximate solution $\tilde{}\bm{x}$. For this article, we seek a vector $\tilde{}\bm{x}$ with a small relative residual:
@@ 1.1 Background and research questions
Most work on matrix–vector complexity for the $\bm{A}\bm{x}=\bm{b}$ problem focuses on the case when $\bm{A}$ is symmetric positive definite (SPD). In this case, the standard Krylov subspace algorithms are conjugate gradient and MINRES [22, Algs. 2 & 4]. The latter achieves the guarantee ˜1.1 using ${\mathcal{O}}(\sqrt{\kappa}\,\log(\nicefrac{{1}}{{\varepsilon}}))$ matvecs. Here, $\kappa\coloneqq\operatorname{cond}(\bm{A})\coloneqq\lVert\bm{A}\rVert_{2}\lVert\smash{\bm{A}^{-1}}\rVert_{2}=\sigma_{\mathrm{max}}(\bm{A})/\sigma_{\mathrm{min}}(\bm{A})$ is the condition number. Classical lower bounds confirm that $\Omega(\sqrt{\kappa}\,\log(\nicefrac{{1}}{{\varepsilon}}))$ matvecs are necessary for any deterministic algorithm to achieve this guarantee [38, Sec. 7.2]. However, with the rapidly increasing use of randomization in computational linear algebra [35, 29, 14], a lower bound against deterministic algorithms is insufficient to understand the difficulty of solving a given problem. This limitation was partially addressed by Braverman et al. [5], who proved a weaker $\Omega(\sqrt{\kappa}/\operatorname{polylog}(\kappa))$ lower bound for SPD linear systems against randomized algorithms.
It is natural to inquire how the matrix–vector complexity (of both deterministic and randomized algorithms) changes when we consider general linear systems for which $\bm{A}$ is not SPD. In this setting, we can distinguish between a two-sided matrix–vector algorithm, which can perform matrix–vector products with both $\bm{A}$ and $\bm{A}^{\smash{\intercal}}$, and a one-sided matrix–vector algorithm, which can only perform matrix–vector products with $\bm{A}$. Both one- and two-sided algorithms are studied and used in practice [22].
@@
The matrix $\bm{A}^{\smash{\intercal}}\bm{A}$ is SPD, so the conjugate gradient method can solve this system. The condition number of the normal equations is larger, $\operatorname{cond}(\bm{A}^{\smash{\intercal}}\bm{A})=\kappa^{2}$, implying that conjugate gradient on the normal equations requires ${\mathcal{O}}(\sqrt{\kappa^{2}}\kern 0.50003pt\log(\nicefrac{{1}}{{\varepsilon}}))={\mathcal{O}}(\kappa\log(\nicefrac{{1}}{{\varepsilon}}))$ matvecs to produce an $\varepsilon$-accurate solution [37, Eqn. 2.12]. This is known to be tight for deterministic algorithms [48, p. 180]. For, randomized algorithms, no existing lower bounds extend beyond the SPD case. We ask:
@@
The most popular one-sided matrix–vector algorithm is the GMRES algorithm [46]. This algorithm has many desirable properties: When it is applied to an SPD system, it achieves the same ${\mathcal{O}}(\sqrt{\kappa}\kern 0.80002pt\log(\nicefrac{{1}}{{\varepsilon}}))$ complexity as MINRES, and it converges rapidly on many problems in practice. However, this algorithm also has a fundamental weakness: There exists a perfectly conditioned linear system ($\kappa=1$) on which GMRES requires a full $n$ iterations to solve the system to any accuracy level $\varepsilon<1$ [37, Ex. C]. This raises the question:
@@ 1.2 Our results
Fix $\eta>0$. There does not exist any randomized algorithm that takes inputs $\kappa$, $\varepsilon$, and $\bm{b}$, which computes fewer than $\frac{1-\eta}{4}\kappa\log(\nicefrac{{1}}{{\varepsilon}})$ two-sided matrix–vector products with $\bm{A}$, and which returns a vector $\tilde{}\bm{x}$ such that
@@
This result shows that conjugate gradient on the normal equations is optimal from the perspective of worst-case complexity. [^2] We note that there are reasons in practice to avoid iterative methods based on the normal equations, including numerical stability [39, Sec. 8.2] and better-than-worst case performance [37]. A strength of this result is its explicitness, with the complexity bound featuring no hidden constants. Conjugate gradient on the normal equations achieves the accuracy guarantee ˜1.1 in $(1+o(1))\kappa\log(\nicefrac{{1}}{{\varepsilon}})$ matvecs [22, Eqn. 3.6], so the lower and upper bounds match to a factor of roughly 4.
@@
Consider any randomized algorithm that takes a vector $\bm{b}$ and a one-sided matrix–vector oracle for a matrix $\bm{A}$ and outputs a vector $\tilde{}\bm{x}$, and suppose $n$ is larger than some universal constant. Then:
$\text{Achieving }\ \lVert\bm{A}\tilde{}\bm{x}-\bm{b}\rVert_{2}$ $\leq\frac{1}{2}\lVert\bm{b}\rVert_{2}$ $\text{with prob.\ }\geq\frac{2}{3}$ $\text{requires at least $\lceil n/2\rceil$ matvecs};$
@@
Algorithms like conjugate gradient and GMRES are already known to be optimal among certain classes of algorithms, like within the class of deterministic single-vector Krylov methods (see [22, Ch. 2] and [38, Sec. 7.2]). However, prior works do not fully answer if these methods are optimal within the larger class of general randomized matrix–vector algorithms. [^3] Though Braverman et al. [5] do prove that $\Omega(\sqrt{\kappa}/\operatorname{polylog}(\kappa))$ matvecs are needed to solve a SPD linear system to constant error, partially resolving this question. Indeed, there have been many examples of randomized algorithms that achieve surprising improvements to the complexity of problems in linear algebra [43, 10, 11, 36, 32, 40, 17]. However, despite all efforts, no improvements to the worst-case complexity of solving general linear systems have materialized. Our research demonstrates why this is the case: Linear system solving is not amenable to such a speedup, in both the one- and two-sided cases.
@@
We use different techniques to handle the one- and two-sided cases. The two-sided lower bound (Theorem˜1.1) uses an approach based on the nonexistence of good polynomial approximations, extending an argument of Chewi et al. [9]. We show that since no polynomial can approximate the function $f(x)=\nicefrac{{1}}{{x}}$ to error $\varepsilon$ on the split interval ${\mathcal{S}}\;{\vcentcolon=}\;[-\kappa,-1]\cup[1,\kappa]$ with degree less than $t={\mathcal{O}}(\kappa\log(1/\varepsilon))$, no algorithm can use fewer than ${\mathcal{O}}(t)$ matvecs to estimate $\operatorname{tr}(f(\bm{A}))=\operatorname{tr}(\bm{A}^{-1})$ for all matrices with eigenvalues in $\mathcal{S}$}. We then make a reduction, showing that any linear system solver satisfying ˜1.1 can be used to estimate $\operatorname{tr}(\bm{A}^{-1})$, resulting in a proof of Theorem˜1.1. See Section˜2 for details.
@@
1.3 Extension: Stronger lower bounds for solving SPD systems
The focus of this article is on solving general, not necessarily SPD, linear systems. Indeed, the prior work of Braverman et al. [5] has already shown that $\Omega(\sqrt{\kappa}/\operatorname{polylog}(\kappa))$ matvecs are needed for a randomized algorithm to solve a SPD linear system. However, our analysis is able to provide tighter lower bounds for the SPD case as well, removing unnecessary polylogarithmic factors in the condition number $\kappa$ and obtain the correct logarithmic dependence on the inverse-accuracy $\nicefrac{{1}}{{\varepsilon}}$:
Theorem 1.3 (SPD Linear systems: Lower bound against two-sided algorithms).
Fix $\eta>0$. There does not exist any randomized algorithm that takes inputs $\kappa$, $\varepsilon$, and $\bm{b}$, which computes fewer than $\frac{1-\eta}{4}\sqrt{\kappa}\log(\nicefrac{{1}}{{\varepsilon}})$ two-sided matrix-vector products with $\bm{A}$, and which returns a vector $\tilde{}\bm{x}$ such that
for all numbers $\varepsilon>0$ and $\kappa\geq 1$ and SPD matrices $\bm{A}$ of condition number $\operatorname{cond}(\bm{A})\leq\kappa$.
This result sharpens the log factors over the prior work and, since conjugate gradient converges in ${\mathcal{O}}(\sqrt{\kappa}\log(\nicefrac{{1}}{{\varepsilon}}))$ matvecs, shows that the dependence on $\varepsilon$ in the upper bound analysis is tight. The proof of Theorem˜1.3 uses the same basic argument as the proof of the lower bound in the general case Theorem˜1.1. See Section˜E.2 for details.
@@ 1.4 Implication: Fine-grained lower bounds and complexity separation
An important and well-studied example of this arises when $\bm{A}$ has a number of large outlying singular values that blow up its condition number, which arises in smoothed analysis [47], optimization [7], and machine learning [50]. In this setting, conjugate gradient on the normal equations is known to solve such linear systems in an improved ${\mathcal{O}}(k+\kappa_{k}\log(1/\varepsilon))$ matvecs [2], where $\kappa_{k}\coloneqq\sigma_{k+1}(\bm{A})/\sigma_{\rm min}(\bm{A})$ denotes the condition number without the top $k$ singular values. This more fine-grained notation of complexity, parameterized by both $k$ and $\kappa_{k}$, circumvents our basic two-sided lower bound (Theorem˜1.1). However, our results can be extended to this different complexity measure, again showing that conjugate gradient on the normal equations is optimal:
@@
There does not exist a randomized algorithm that takes inputs $k$, $\kappa_{k}$, $\varepsilon$, and $\bm{b}$, which computes fewer than $\Omega(k+\kappa_{k}\log(1/\varepsilon))$ two-sided matrix-vector products with $\bm{A}$ and returns $\tilde{}\bm{x}$ such that
@@
Notably, this result implies a complexity separation between matrix–vector algorithms and general algorithms that can access the entries of $\bm{A}$ directly. If $\bm{A}$ is a dense matrix and we implement matvecs using the standard algorithm, then Theorem˜1.4 shows that $\Omega(n^{2}k+n^{2}\kappa_{k}\log(1/\varepsilon))$ time is needed to solve a linear system. However, recent work [19, 16, 17] shows that if $\bm{A}$ is stored as a dense array, we can solve $\bm{A}\bm{x}=\bm{b}$ in $\widetilde{}{\mathcal{O}}(k^{3}+n^{2}\kappa_{k}\log(1/\varepsilon))$ time. So, for large $n$ and $k$, having access to the matrix offers a direct improvement over matrix–vector access, demonstrating a complexity separation between matrix–vector algorithms and general algorithms. This is the first result of this kind for general, dense linear systems.
@@ 1.5 Related work
Understanding the behavior of Krylov linear solvers has been a decades-long research challenge in numerical analysis. The literature is extensive; see Greenbaum’s book [22] for a summary of the theory. Lower bounds for the rate of convergence of Krylov methods can be derived from potential theory [18]. These lower bounds apply only to specific Krylov methods, not general randomized matrix–vector algorithms.
@@
Our work fits into the growing research topic of transpose-free linear algebra [4, 23], which aims to understand the relative power of one- and two-sided algorithms for various linear algebra problems. The results most similar to ours appear in the thesis of Halikias [23, Sec. 5.2], who analyzes one-sided matrix–vector algorithms for solving linear systems and least-squares problems. Her results show that, even in the favorable situation when $\operatorname{cond}(\bm{A})=1$ and we perform n1n-1 matvecs $\bm{A}\bm{z}_{1},\ldots,\bm{A}\bm{z}_{n-1}$, [^4] Halikias also includes as an assumption that the exact solution $\bm{A}^{-1}\bm{b}$ is not in the span of the queries $\{\bm{z}_{1},\ldots,\bm{z}_{n-1}\}$. the data collected by the algorithm is not sufficient to characterize the solution to a system of linear equations: There exists another well-conditioned matrix $\tilde{}\bm{A}$ that is undistinguished by the queries
@@
Our $\Omega(\kappa)$ lower bound for two-sided matrix–vector algorithms mirrors the $\Omega(\kappa)$ lower bound for the quantum linear systems problem [24, p. 150502-3]. In that context, solving implicit sparse systems in ${\mathcal{O}}(\kappa^{1-\eta}\operatorname{polylog}(n))$ time would imply $\mathsf{BQP}=\mathsf{PSPACE}$ and allow for sub-linear simulation of arbitrary quantum circuits. While the $\Omega(\kappa)$ scaling in both settings is striking, it remains unclear how to derive one result from the other or reduce them both to some common principle.